OpenStudy (anonymous):

what is the equation for the arithmentic sequence where the first number is 0 and the pattern increases by 11 each time?

OpenStudy (anonymous):

is it an=0+11(n-1)

OpenStudy (xapproachesinfinity):

well the first term is 0 then the second term will be 11, third term is 22, fourth term is 33 and so forth since the common difference is 11 An=A+(n-1)D where A is the first term A=0 in this case D is the common difference that you are adding each time or subtracting here you D=11

OpenStudy (anonymous):

so I was correct?

OpenStudy (xapproachesinfinity):

construct is like this A=0 A2=0+(2-1)11 A3=0+(3-1)11 A4=0+(4-1)11 . . . . . . An=A1+(n-1)11

OpenStudy (xapproachesinfinity):

yes!

OpenStudy (anonymous):

thank you! @xapproachesinfinity

OpenStudy (xapproachesinfinity):

welcome! even if you don't know what is the general term you can always construct to it using what i said above start with the first term and build^_^